OpenStudy (anonymous):

The region bounded by the curve y=sqrt(x) and the lines x=0 and y=6 rotated about the line x=-2

OpenStudy (tkhunny):

What have you tried? What method do you prefer?

OpenStudy (anonymous):

Washer method

OpenStudy (tkhunny):

Okay, then set it up: \(\pi \int r^{2}\;dh\) Go!

OpenStudy (anonymous):

Isnt there an inner and outer radius?

OpenStudy (tkhunny):

There is, but the inner radius is zero (0).

OpenStudy (tkhunny):

Oh, sorry, I read x = 0 and x = 6. Rethink and find the inner radius. x = 0 and y = 6!

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

You will need the intersection of y = 6 and y = sqrt(x).

OpenStudy (anonymous):

ya (6,36)

OpenStudy (tkhunny):

Okay, the Outer Radius is a constant 8. What is the Inner Radius?

OpenStudy (anonymous):

should be 2

OpenStudy (tkhunny):

It is 2 ONLY for x = 0. How about for the rest of the Domain, (0,36]?

OpenStudy (anonymous):

i have no clue

OpenStudy (tkhunny):

The you cannot solve this problem. Give it a go. What is the inner radius for x = 1? \(\sqrt{1} = 1\), so the Inner Radius there would be 1-(-2) = 1+2 = 3 How about x = 4?

OpenStudy (anonymous):

4

OpenStudy (tkhunny):

How did you get that?

OpenStudy (anonymous):

does 8-y^2 have something to do with it

OpenStudy (tkhunny):

You didn't answer by question and no.

OpenStudy (anonymous):

i literally just plugged in 4 for x

OpenStudy (tkhunny):

Oh, so you used the actual function to determine the radius. That's it. Your inner radius is \(\sqrt{x}\). Okay, now the limits.

OpenStudy (tkhunny):

Oh, so you used the actual function to determine the radius. That's it. Your inner radius is \(\sqrt{x}+2\). Okay, now the limits.

OpenStudy (anonymous):

one should be 36

OpenStudy (anonymous):

oops, 6

OpenStudy (anonymous):

the other maybe -2

OpenStudy (tkhunny):

The radius is parallel to the y-axis. Your limits need to be on the x-axis, perpendicular to your radius.

OpenStudy (anonymous):

so 0 and 36

OpenStudy (tkhunny):

Done. What do you get?

OpenStudy (anonymous):

i got 936 but i know thats not right

OpenStudy (tkhunny):

You missed the \(\pi\).

OpenStudy (tkhunny):

Alternatively, with Shells \(2\pi\int\limits_{0}^{6}(y+2)\cdot y^{2}\;dy\) Same thing.

OpenStudy (anonymous):

u sure that is right?

OpenStudy (anonymous):

the 936pi

OpenStudy (tkhunny):

That's why I like to do it both ways. If I don't get the same answer both ways, I keep working at it. Yes!

OpenStudy (anonymous):

the outer radius was 8 and inner sqrt(x)+2 correct?

OpenStudy (tkhunny):

Indeed.

OpenStudy (tkhunny):

\(\pi\int\limits_{0}^{36}8^{2} - (\sqrt{x}+2)^{2}\;dx\)

OpenStudy (anonymous):

my online thing says its wrong

OpenStudy (anonymous):

oh well thanks brother

OpenStudy (tkhunny):

It's possible we don't have the right problem statement. This is the solution to the problem presented.

OpenStudy (anonymous):

the region bounded by the curve y=sqrt(x) and the lines x=0 and y=6. Then it asks me to find the volume of the solid when the region is rotated about the line x= -2

OpenStudy (tkhunny):

I believe you. That is the problem we solved. I cannot speak for the online acceptance. There is an error in this process, somewhere. It is not in the solution to the problem presented. Does it want a purely numerical response? The numbers in this problem are rather large. There may be some rounding problem. Try 2940.5290987556 and 2940.5307237601

OpenStudy (anonymous):

ill ask my math teacher about it tuesday, i LOVE high school ap calc

OpenStudy (tkhunny):

Go get 'em. The more I see online attempts at math education, without real humans, the less I am impressed. Maybe one day we'll start to get better at it.

OpenStudy (anonymous):

true that man, true that